If a boy can throw a stone of maximum height of 25m, he can throw the same stone to a distance of ?

Grade:8

1 Answers

Arun
6 years ago
 
Given vertical height = 25m. if we took it as max vertical height
then, u2/2g = 25m => u2/g = 50
 
As horizontal projectile range = u2sin2∂/g
So, range is maximum when ∂=45 degrees
 
Hence, maximum horizontal range is u2/g = 50,m